### 12 Mar 2026 v1

# UtilityMax Prompting: A Formal Framework for Multi-Objective Large Language Model Optimization

Abstract
The success of a Large Language Model (LLM) task depends heavily on its prompt. Most use-cases specify prompts using natural language, which is inherently ambiguous when multiple objectives must be simultaneously satisfied. In this paper we introduce UtilityMax Prompting, a framework that specifies tasks using formal mathematical language. We reconstruct the task as an influence diagram in which the LLM’s answer is the sole decision variable. A utility function is defined over the conditional probability distributions within the diagram, and the LLM is instructed to find the answer that maximises expected utility. This constrains the LLM to reason explicitly about each component of the objective, directing its output toward a precise optimization target rather than a subjective natural language interpretation. We validate our approach on the MovieLens 1M dataset across three frontier models (Claude Sonnet 4.6, GPT-5.4, and Gemini 2.5 Pro), demonstrating consistent improvements in precision and Normalized Discounted Cumulative Gain (NDCG) over natural language baselines in a multi-objective movie recommendation task.
1 Introduction
Prompt engineering is crucial in directing Large Language Models (LLMs) to solve specific tasks. Various prompt engineering techniques have been proposed in recent years Sahoo et al. (2024). In particular, zero-shot prompting is attractive due to its simplicity since it does not require external exemplars to construct the prompt Li (2023); Radford et al. (2019).
One line of research focuses on restructuring how the model reasons within a single inference pass. With zero-shot Chain-of-Thought (CoT) prompting, the LLM is instructed to reason “step-by-step” about the solution before producing an answer Wei et al. (2022); Cheng et al. (2025). In the few-shot CoT setting, exemplars of similar questions and answers are provided so that the LLM can observe the reasoning patterns it should follow Wei et al. (2022). In Program of Thoughts (PoT) prompting, the LLM translates a natural language prompt into a programming language such as Python, which is then executed to produce the final answer Chen et al. (2023). In Chain-of-Symbol (CoS) prompting, a task is converted into a symbolic representation that assists the LLM in solving spatial reasoning and planning tasks Hu et al. (2023).
A second line of research treats the prompt itself as the optimization target. For example, Optimization by Prompting (OPRO) uses an LLM as an iterative optimizer, generating candidate prompts, evaluating them against a labeled dataset, and refining them across multiple passes Yang et al. (2024). While effective, this approach requires a scoring function for the task so that proposed answers can be evaluated as better or worse. Such evaluation signals are not always available, or are expensive to obtain in many real-world deployment settings.
These methods have demonstrated significant improvements in LLM performance across a range of benchmarks. However, they share a common assumption: the task objective is specified in natural language. While this is not a limitation in domains with a single objective, such as solving mathematical problems, it becomes a challenge when multiple dependent objectives must be optimized simultaneously.
For example, consider a trading agent that wants to maximise profit at a given level of risk tolerance. These objectives naturally compete with each other as the most profitable strategy may require taking a level of risk that could result in significant losses. Balancing both objectives therefore becomes key, but a natural language prompt that says: “maximise profit subject to a medium level of risk” is inherently ambiguous. The LLM has to interpret what “medium” means here. While a more carefully crafted natural language prompt could partially resolve this kind of ambiguity, a task like this is best specified in formal mathematical language that eliminates ambiguity entirely.
In this paper we introduce UtilityMax Prompting, a zero-shot framework that takes a complementary approach: rather than restructuring the reasoning process or iterating over candidate prompts, we replace the natural language objective itself with a formal mathematical specification. The LLM is instructed to maximise the expected utility given its answer, constraining it to reason about each component of the objective individually and directing its output toward a well-defined optimization target. This approach requires neither exemplars nor a scoring function.
The rest of this paper is organized as follows: in Section 2 we cover the general methodology behind the UtilityMax framework; in Section 3 we extend the framework to handle dependencies between binary chance nodes; in Section 4 we provide a prompting template under UtilityMax that practitioners can use; in Section 5 we validate UtilityMax on a multi-objective movie recommendation task using the MovieLens 1M dataset Harper and Konstan (2015), demonstrating consistent improvements in precision and Normalized Discounted Cumulative Gain (NDCG) over natural language baselines across three frontier models (Claude Sonnet 4.6, GPT-5.4, and Gemini 2.5 Pro); and we conclude with final remarks and directions for future research in Section 6.
2 Framework
Let $K$ represent the LLM’s knowledge. This includes all internal knowledge stored through its parameters as well as external knowledge such as context through chat history or online research. Given $K$, we define the following influence diagram:
Let $A\mid K$ be a decision node that represents the space of all possible LLM’s answers given $K$. An answer $a\in A$ is decided by the LLM. Let $\{X_{1},X_{2},X_{3},...,X_{n}\}$ be a set of chance nodes with $n\geq 1$. We require:
  1. 1.
    The dependency structure over $\{A,X_{1},...,X_{n}\}$ forms a directed acyclic graph (DAG) with $A$ as the singular root node.
  2. 2.
    Each $X_{i}$ is conditionally independent given $A$.
Suppose further that we define a multiplicative utility function $U(X_{1},...,X_{n})=\prod_{i=1}^{n}f_{i}(X_{i})$. Then the expected utility factorizes as:
$\mathbb{E}[U\mid A]=\mathbb{E}\!\left[\prod_{i=1}^{n}f_{i}(X_{i})\mid A\right]=\prod_{i=1}^{n}\mathbb{E}[f_{i}(X_{i})\mid A].$ (1)
This setup is shown visually in Figure 1.
[…]
Figure 1: Influence diagram for UtilityMax Prompting showing decision node AA, chance nodes $\{X_{1},X_{2},X_{3},...,X_{n}\}$, and utility node $U$.
Then the task of the LLM is to find $a^{*}\in A$ that maximises $\mathbb{E}[U\mid A=a^{*}]$.
3 Binary Chance Nodes
In the special case where the chance nodes $\{X_{1},X_{2},\ldots,X_{n}\}$ are all binary random variables, we can relax the conditional independence assumption. Let $pa(X_{i})$ denote the parents of $X_{i}$ in the DAG, and suppose that if any $X_{j}=0$ for $X_{j}\in pa(X_{i})$ then $P(X_{i}=1\mid pa(X_{i}),A)=0$. That is, each node is deterministically gated by its parents so that a child node can only be active if all its parents are active.
Denote by $pa(X_{i})=1$ the event that $X_{j}=1$ for all $X_{j}\in pa(X_{i})$. Under the gating assumption, all terms in the expectation where any parent takes the value zero vanish, and the expected utility reduces to:
$\mathbb{E}[U\mid A]=\prod_{i=1}^{n}P(X_{i}=1\mid pa(X_{i})=1,A)$ (2)
where we define $f_{i}(X_{i})=X_{i}$ so that $\mathbb{E}[X_{i}\mid pa(X_{i})=1,A]=P(X_{i}=1\mid pa(X_{i})=1,A)$. Each term is estimated independently by the LLM, preserving tractability without requiring conditional independence.
4 Prompting Template
Given the framework outlined in Section 2, we can construct a prompt to maximise the given utility. A template for such a prompt is given below for two random variables $X_{1}$ and $X_{2}$ .
I want you to solve the following task: [TASK DESCRIPTION].
Formally, let K represent your knowledge. This includes all your internal knowledge stored through your parameters as well as any external knowledge provided in this prompt or chat history.
Let P(A | K) represent your probability distribution over answers given K. Let a be an answer in A.
Let X1 | A=a be a random variable representing [DESCRIPTION OF X1] given answer a.
Let X2 | A=a be a random variable representing [DESCRIPTION OF X2] given answer a.
Your task is to use your domain expertise to find the optimal answer a* that maximises O(a) = E[X1 | A=a] x E[X2 | A=a]. To do this you must:
1. Generate a set of candidate answers.
2. For each candidate answer, estimate E[X1 | A=a] and E[X2 | A=a] individually using your internal knowledge then compute O(a) for that candidate.
3. Return the answer a* that maximises O.
5 Experimental Setup
To validate our prompting framework we run experiments on the MovieLens 1M dataset Harper and Konstan (2015). Given the first $100$ movies that a user has rated between $1$ and $5$, the LLM’s task is to recommend the top $10$ movies for that user from the next $50$ in their watched list.
To cast the problem as a multi-objective task, we further condition that the user is only interested in movies in the comedy and romance genres. To ensure sufficient signal in the test set, we only select users that have at least $5$ movies in the test window that belong to both genres and carry a rating of $4$ or higher, and that have a total watch history of at least $150$ movies. We randomly select $20$ users and for each user we run $20$ queries to account for stochasticity in LLM outputs. We evaluate three prompt types:
  1. 1.
    Basic: The LLM is told that the user is in the mood for comedy and romance movies.
  2. 2.
    Harsh: The LLM is told that the user is only interested in comedy and romance movies and that it should not suggest anything outside of these genres.
  3. 3.
    UtilityMax: A prompt is constructed based on the template of Section 4 with three random variables - a categorical random variable $S$ for the predicted score, a binary random variable $G_{1}$ for the comedy genre, and a binary random variable $G_{2}$ for the romance genre. Therefore, the objective is:
    $O(a)=\mathbb{E}[S\mid A=a]\times P(G_{1}=1\mid A=a)\times P(G_{2}=1\mid A=a)$ (3)
The LLM is not given the test movie scores, nor the genre labels for any movie in either the training or test set. Under UtilityMax, the LLM must therefore estimate $\mathbb{E}[S\mid A=a]$, $P(G_{1}=1\mid A=a)$, and $P(G_{2}=1\mid A=a)$ for each candidate movie using only its title and the user’s training history.
The recommended list is evaluated against the test set, where a movie is considered a positive match only if it carries a rating of $4$ or higher and belongs to both the comedy and romance genres. We report Precision@10 and Normalised Discounted Cumulative Gain (NDCG@10) averaged across all users and runs. Experiments are conducted across three frontier models: Claude Sonnet 4.6, GPT-5.4, and Gemini 2.5 Pro.
Our results are summarized in Table 1. UtilityMax outperforms both Basic and Harsh prompt types across all three models on both metrics. For example, with Claude Sonnet 4.6 we observe a $12.7\%$ and $16.5\%$ improvement in Precision@10 and NDCG@10 respectively over the Basic prompt, and a $11.9\%$ and $18.8\%$ improvement over the Harsh prompt.
Model Prompt Precision@10 NDCG@10
Claude Sonnet 4.6 Basic 0.418 (0.054) 0.570 (0.070)
Harsh 0.421 (0.067) 0.559 (0.083)
UtilityMax 0.471 (0.060) 0.664 (0.064)
GPT-5.4 Basic 0.532 (0.052) 0.739 (0.056)
Harsh 0.518 (0.050) 0.712 (0.060)
UtilityMax 0.578 (0.047) 0.788 (0.050)
Gemini 2.5 Pro Basic 0.417 (0.073) 0.575 (0.098)
Harsh 0.449 (0.086) 0.601 (0.108)
UtilityMax 0.496 (0.079) 0.667 (0.103)
Table 1: Precision@10 and NDCG@10 averaged over 2020 users and 2020 queries per user. Standard deviations are shown in parentheses.
It is interesting to note that the Harsh prompt type does not consistently outperform Basic across models. Harsh underperforms Basic on NDCG@10 for Claude Sonnet 4.6 (0.559 vs 0.570) and on both metrics for GPT-5.4. This inconsistency suggests that increasing the forcefulness of a natural language objective does not reliably resolve the underlying ambiguity in how multiple objectives are weighted. UtilityMax eliminates this ambiguity entirely through formal mathematical specification, which we believe accounts for its consistent superiority across all models and metrics.
GPT-5.4 achieves substantially higher absolute scores across all prompt types compared to the other two models. This may reflect training data overlap between GPT-5.4 and the MovieLens dataset, which is widely used and publicly available. However, the key finding for this paper is that UtilityMax continues to outperform both natural language baselines even for GPT-5.4, suggesting that the formal objective provides genuine additional signal regardless of model capability.
To assess statistical significance we apply a one-sided paired Wilcoxon signed-rank test on per-user mean NDCG@10, comparing UtilityMax against each baseline independently. The null hypothesis is that UtilityMax and the baseline produce equal NDCG scores; the alternative is that UtilityMax is superior. As shown in Table 2, UtilityMax significantly outperforms both baselines across all three models ($p<0.01$ in all cases).
Model vs Basic vs Harsh
Claude Sonnet 4.6 0.0028 0.0006
GPT-5.4 0.0068 0.0042
Gemini 2.5 Pro 0.0014 0.0032
Table 2: One-sided paired Wilcoxon signed-rank test p-values for UtilityMax vs each baseline on NDCG@10. All results are significant at the $p<0.01$ level.
Finally, we note that the effectiveness of UtilityMax depends on the underlying model’s ability to produce well-calibrated probability estimates. A model that cannot reliably estimate the objective components may not benefit from this framework, and may even perform worse than natural language prompting. The results presented here suggest that current frontier models are capable enough to leverage the formal objective effectively, but we anticipate that weaker models may fall below this capability threshold. We leave a systematic investigation of this threshold to future work.
6 Final Remarks
This paper has introduced UtilityMax, a formal zero-shot prompting framework for multi-objective LLM optimization. The key insight is that by representing the task as a formal optimization objective, the LLM is constrained to reason explicitly about each component of that objective, rather than relying on a natural language interpretation that may be ambiguous.
The effectiveness of UtilityMax depends on two key factors. First, the underlying model must be expressive enough to produce well-calibrated probability estimates for the objective components. Our experiments suggest that current frontier models meet this requirement, at least on the MovieLens recommendation task.
Second, the optimization objective must be carefully designed so that its variables capture the most important components of the underlying task. Including irrelevant variables introduces redundancy and increases computational cost, while omitting important variables will cause the framework to optimize a proxy that does not fully reflect the true goal. The designer must therefore exercise judgement in selecting variables that are both necessary and sufficient for the task at hand.
Several directions for future research present themselves. Perhaps the most practically important is automating the construction of the UtilityMax prompt. Specifically, developing a method by which an LLM can extract the relevant variables from a natural language task description and format them into the UtilityMax framework automatically. This would remove the requirement for a human designer to manually specify the optimization objective.
A second important direction is further relaxing the conditional independence assumption on the chance nodes given $A$. While this assumption yields a tractable objective and is sufficient to capture a wide range of tasks, it precludes dependencies between chance nodes that may be necessary in certain domains. Extending the framework beyond the gating mechanism introduced in Section 3 to handle such dependencies while preserving tractability remains an open problem.
Finally, a systematic investigation of the capability threshold below which UtilityMax ceases to be beneficial would help practitioners identify which models are suitable for this framework.
Overall, UtilityMax is a promising approach to improving LLM task performance through formal objective specification. Our experimental evidence across three frontier models demonstrates consistent improvements over natural language baselines, and we hope this work motivates further research into the formal specification of LLM objectives as a complement to existing prompt engineering techniques.